7. a) At a call centre, callers have to wait till an operator is ready to take their call. To

monitor this process, 5 calls were recorded every hour for the 8-hour working day.

The data below shows the waiting time in weeks:

Time Sample number

1 2 3 4 5

9 a.m. 8 9 15 4 11

10 7 10 7 6 8

11 11 12 10 9 10

12 12 8 6 9 12

1 p.m. 11 10 6 14 11

2 7 7 10 4 11

3 10 7 4 10 10

4 8 11 11 11 7

Compute the control limits for x and R -charts. Draw the control charts for x and

,R and comments.


1
Expert's answer
2021-05-11T09:08:21-0400

Given "\mu_0=26, \bar{x}=20, s=10, n=18."


The following null and alternative hypotheses need to be tested:

"H_0:\mu=26"

"H_1:\mu\not=26"


This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.


Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test with degrees of freedom "df=n-1=18-1=17" is "t_c=2.109816."


The rejection region for this two-tailed test is "R=\{t:|t|>2.109816\}."

  

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu_0}{s/\sqrt{n}}=\dfrac{20-26}{10/\sqrt{18}}=-2.545584"


Since it is observed that "|t|=2.545584>2.109816=t_c," it is then concluded that the null hypothesis is rejected.


Therefore, there is enough evidence to claim that the population mean "\mu" is different than "26\ cm," at the "0.05" significance level.


Using the P-value approach:


The p-value for two-tailed, "df=17, \alpha=0.05, t=-2.545584" is "p=0.020881," and since "p=0.020881<0.05=\alpha," it is then concluded that the null hypothesis is rejected.


Therefore, there is enough evidence to claim that the population mean "\mu" is different than "26\ cm," at the "0.05" significance level.


Range mean:


"\bar{R}=\frac{\sum R}{n}"


"\bar{R}=\frac{9.66}{25}"


"\bar{R}=0.386"


Control limits for mean chart:


"UCL_{\bar{x}} = \bar{\bar{x}} +A_{2}\bar{R}"


 = 20.677 + (0.419)(0.386)


 = 20.83


"CL_{\bar{x}} =\bar{\bar{x}} = 20.677"


"LCL_{\bar{x}} =\bar{\bar{x}} - A_{2}\bar{R}"


"LCL_{\bar{x}} =20.677-(0.419)(0.386)"


"LCL_{\bar{x}} =20.51"


Control limits for range chart:


"LCL_{\bar{R}} =D_{3}\bar{R} = 0.076*0.386 = 0.029"


"CL_{\bar{R}} =0.386"


"UCL_{\bar{R}} =D_{4}\bar{R} = 1.924\times 0.386 = 0.74"


Mean chart is-




Range chart-